Services operational

About This Site

We'll report any known service issues on this page, including any planned maintenance.

You can contact us via your customer portal, our contact page or via phone on 0333 773 3000.

Past Incidents

Thursday 22nd July 2021

Shared Hosting november - Non-Responsive

We are investigating non-response of the november server.

The server has remained stable and we consider this issue now resolved.

Thank you for your understanding and patience throughout.

The server has been running normally now for 1 hour.

It appears there may have been a power surge in the datacentre as an unrelated network device and remote access tool have also both failed. On inspecting all three devices they all show signs of PSU (power supply unit) failures.

We will take this up with the datacentre.

We need to reboot once due to a clock sync issue. This should not take longer than a few minutes.

Whilst restoring to new hardware, we have been able to boot up the previous images.

The server is now online and we will look in to the failed hardware in more detail to find the root case.

We are still working on the restoration of data of this server.

Thank you for your patience.

The card and drives have been moved and the replacement chassis is booting.

Unfortunately, it appears not to want to boot into the array. We are attempting some further things but should that fail, our next step will be to restore from backup.

The engineer is moving the RAID card and drives to an identical chassis and if the new chassis boots, we will then shutdown to re-add the original memory.

In cases where this works normally, either the motherboard or CPU would have failed in the original chassis.

Such failures are rare. In over 15 years, we've only seen three servers have serious hardware failures but we appreciate the impact it will be having right now which we apologise for.

The hardware on this server appears to have failed but the exact cause cannot be isolated. We are seeing power to the server but it is not powering up.

A chassis swap is being performed in an attempt at the quickest route to restore service but should this fail, it will be a restoration from the last backup overnight to a new server.

We will provide an update as soon as we can and appreciate your patience.

The engineer is still working on this server.

Initial findings indicate a potential hardware failure but checks are ongoing.

An engineer is working on the server at the datacentre and we will provide an update as quick as possible.

Wednesday 21st July 2021

No incidents reported